A buffer solution with a pH of 4.25 is prepared with the same volumes of _________ M formic acid and 0.12 M sodium formate.

The Ka of formic acid is 1.8 ⋅ 10-4.

3.7×10−2

0.38

1.8×104

7.5×10−2

1.9×10−2
Answer #1

pH = 4.25

Ka = 1.8*10-4

pKa = -log Ka = -log (1.8*10-4) = 3.74

[HCOO-] = 0.12 M

[HCOOH] = ?

Using Handerson Hasselback equation:

pH = pKa + log [HCOO-]/[HCOOH]

4.25 = 3.74 + log (0.12)/[HCOOH]

log [(0.12)/[HCOOH]] = 0.51

[(0.12)/[HCOOH]] = 3.24

[HCOOH] = 0.12/3.24 = 0.037

[HCOOH] = 0.037 M = 3.7*10-2 M

1st option is correct